# Event-based Sensor Fusion and Application on Odometry: A Survey

### v2

@@ Abstract
Event cameras, inspired by biological vision, are asynchronous sensors that detect changes in brightness, offering notable advantages in environments characterized by high-speed motion, low lighting, or wide dynamic range. These distinctive properties render event cameras particularly effective for sensor fusion in robotics and computer vision, especially in enhancing traditional visual or LiDAR-inertial odometry. Conventional frame-based cameras suffer from limitations such as motion blur and drift, which can be mitigated by the continuous, low-latency data provided by event cameras. Similarly, LiDAR-based odometry encounters challenges related to the loss of geometric information in environments such as corridors. To address these limitations, unlike the existing event camera-related surveys, this paper presents a comprehensive overview of recent advancements in event-based sensor fusion for odometry applications particularly, investigating fusion strategies that incorporate frame-based cameras, inertial measurement units (IMUs), and LiDAR. The survey critically assesses the contributions of these fusion methods to improving odometry performance in complex environments, while highlighting key applications, and discussing the strengths, limitations, and unresolved challenges. Additionally, it offers insights into potential future research directions to advance event-based sensor fusion for next-generation odometry applications.
@@ I Introduction
In the context of robotics and computer vision, sensor fusion involving event cameras has garnered considerable attention, particularly for visual odometry (VO) or visual-inertial odometry (VIO). Visual odometry aims to estimate a robot’s trajectory by analyzing visual inputs, often combined with other sensor data. Event-based sensor fusion, therefore, seeks to leverage the complementary strengths of event cameras and other sensors (e.g., frame-based cameras, inertial measurement units (IMUs), LiDAR) to improve robustness and accuracy in hash environments with high-speed needs. Regarding odometry, LiDAR-based systems generally offer greater accuracy and robustness to varying lighting conditions compared to visual odometry (VO) or visual-inertial odometry (VIO). However, in environments where geometric information is homogenous, such as long corridors, LiDAR odometry is prone to drift [4]. Event cameras, with their ability to capture dynamic scene changes, have the potential to mitigate this issue by providing complementary information that enhances the overall accuracy and robustness of odometry in such challenging scenarios. Figure 1 illustrates the varying representations of a long corridor as captured by different sensors, including an RGB camera, LiDAR, and an event camera. It is evident from the LiDAR data that the corridor is not clearly detected; however, by integrating the data from the RGB camera and the event camera, this limitation can be mitigated.
@@
The remainder of this paper is structured as follows: Section II provides an overview of the background and fundamentals of event cameras. Following that, Section III discusses the key trends in event camera-based sensor fusion tasks particularly on enhancing odometry by integrating event cameras with other sensors. Section IV introduces the most recent datasets relevant to event cameras for odometry. Finally, Section V concludes the paper, offering insights into potential future research directions.
II Background of the Event Camera
Event cameras capture changes in brightness asynchronously for each pixel. Unlike traditional frame-based cameras that capture intensity images at fixed intervals, event cameras detect changes in the logarithmic brightness of individual pixels independently, producing events whenever a change exceeds a predefined threshold [6]. Specifically, an event at a pixel is triggered if the log intensity changes since the last event surpasses the contrast threshold:
$\Delta L(x,y,t)=\log I(x,y,t)-\log I(x,y,t_{\text{last}})=p\cdot C$ (1)
where $\Delta L(x,y,t)$ represents the logarithmic brightness at the pixel at a time; $I(x,y,t)$ is the intensity at pixel $(x,y)$ at time $t$; $t_{\text{last}}$ is the time of the last event at pixel $(x,y)$; $p\in\{+1,-1\}$ represents the polarity of the event; $C$ is the contrast threshold. This threshold determines the camera’s sensitivity to brightness changes.
Currently, there are multiple primary event camera manufacturers which are listed in Table I alongside several key features illustrated.
@@ II-A Comparison with Traditional Cameras
Event cameras offer several advantages over traditional frame-based cameras, which make them particularly suitable for specific applications in robotics and computer vision [3, 6]. Unlike traditional cameras that capture images at fixed frame rates, event cameras capture changes in brightness asynchronously for each pixel. This difference leads to several key advantages:
@@
II-B Sensor Fusion for Odometry in Robotics
Sensor fusion involves integrating data from multiple sensors to achieve more reliable and accurate measurements. In the context of odometry estimation for robotics and autonomous systems, several key sensors are commonly utilized, including visual sensors, LiDAR, IMU, and others.
Traditional VO methods can suffer from challenges such as motion blur, occlusion, and sensitivity to lighting variations. Event cameras, with their high temporal resolution and ability to capture asynchronous changes, offer a promising solution to address these challenges by providing complementary information that enhances the robustness and accuracy of VO.
To enhance robustness, several studies have explored different sensors like LiDAR, IMU, and radar. Fast-LIO [10][11] combined LiDAR and IMU sensors to achieve odometry, where LiDAR points were extracted as planar and edge features. An iterative Kalman filter was used to fuse these features with IMU measurements, providing accurate state estimation. Building on this, COIN-LIO [12] introduced an improvement by incorporating photometric information from LiDAR intensity data. The system projected LiDAR intensity values into an image format and applied filtering to enhance brightness consistency. The key features in these intensity images were selected based on their complementarity to the geometric features of the LiDAR point cloud, specifically targeting areas with limited geometric information.
@@ II-B Sensor Fusion for Odometry in Robotics
III Event-based Sensor Fusion for Odometry
III-A Event-Camera Only
@@
Since the Event camera only detects the change in brightness, the features that can be used for localization are less than frame cameras. Hence, the amount of studies using only an event camera for odometry is small. The first event-based pose estimation work was presented by Weikersdorfer et al. [16] in 2012. This work used only an event camera and was extended to an event-based visual SLAM system in 2013 [17].
III-B Fusion with Frame-based Camera
@@
Combining event cameras with frame-based cameras allows for high temporal resolution with more structured scene information. In [18], the authors present an approach to combine both sensor modalities for object detection and tracking. In [19], the authors propose a method to reconstruct high frame-rate, sharp videos from a single blurry frame and event data using the Event-based Double Integral (EDI) model. The fusion of asynchronous event data with frame-based data reduces blur and improves visual quality, showcasing the complementary strengths of both modalities. Fusing an event camera with a frame camera can overcome the limitation of each separate sensor and estimate the stereo depth. For example, the work in [20] enables stereo depth estimation by leveraging both event-based edge information and frame-based intensity data, offering improved disparity estimation compared to state-of-the-art stereo matching algorithms.
@@
The fusion of an event camera and a frame camera can increase the robustness and accuracy of visual odometry (VO). In 2014, Andrea Censi and Davide Scaramuzza proposed a low-latency event-based VO system [21], which was a significant advancement in reducing the latency associated with traditional frame-based methods. This system utilized asynchronous event data to compute odometry in real time, paving the way for faster and more efficient pose estimation in dynamic environments. Then in 2016, the work further enhanced the field by introducing low-latency VO using event-based feature tracks [22]. This method tracked feature points based on event data, achieving reliable and low-latency odometry even in conditions where traditional frame-based methods would struggle, such as in fast motion or poor lighting. The same group developed a 6-degree-of-freedom (DoF) event-based tracking and mapping system, which relied on geometric principles to perform both parallel tracking and mapping in real time [23]. This work was instrumental in showcasing the potential of event cameras for performing full 3D tracking and mapping tasks, marking another milestone in event-based VO.
@@
In 2022, Hidalgo-Carrión et al. proposed event-aided direct sparse odometry (DSO), where event data was combined with traditional DSO to improve the robustness of VO in environments with challenging lighting or rapid motion changes [24]. This method demonstrated the fusion of event-based and frame-based approaches in improving VO performance.
III-C Fusion with IMU
@@
IMUs provide inertial data that can be fused with events to enhance motion estimation, particularly in high-speed scenarios. IMUs can provide estimations of velocity and orientation, aiding visual odometry by reducing drift and improving robustness [25]. In [26], the authors explore the use of event data and IMU fusion to achieve enhanced VO performance in dynamic environments. The study proposes methods to leverage the event camera’s high temporal resolution with the IMU’s robust inertial data to accurately estimate motion, even in highly dynamic scenarios Another notable work is [27], the authors propose new methods for integrating event data and IMU readings to achieve ultra-fast camera pose estimates. They introduce two novel fusion schemes that combine constant velocity and constant acceleration prediction models with ultra-fast (10 kHz) event-based updates and slower 1 kHz IMU updates. The findings demonstrate superior performance at a throughput that is 100 times faster than the state-of-the-art.
@@
Integrating an IMU into VO systems significantly enhances their performance by providing additional, reliable motion information that compensates for the limitations of visual data alone. In 2017, Zhu et al. extended the capabilities of event-based odometry by integrating inertial measurements in their work on event-based visual-inertial odometry (EVIO) [28]. The combination of event cameras and inertial sensors allowed for robust pose estimation even in challenging scenarios, such as high-speed motion or complex dynamic scenes. In 2018, Elias Mueggler and his team presented a continuous-time visual-inertial odometry approach for event cameras, further enhancing the accuracy and robustness of event-based odometry systems [29]. Their work highlighted the importance of fusing temporal information from both event data and inertial sensors in a continuous-time framework.
@@
Florian Mahlknecht et al. explored event camera-based odometry for planetary robots in 2022, showcasing the potential of event cameras in extraterrestrial environments where traditional cameras may struggle due to harsh conditions [30]. In 2023, the authors in [31] introduced a hybrid approach combining event and frame-based visual-inertial odometry with adaptive filtering techniques based on 8-DoF warping uncertainty. This method enhanced the odometry system’s resilience to noise and dynamic changes, further improving the robustness and adaptability of event-based systems in various application settings.
III-D Event-based Stereo Visual Odometry
@@
Event-based stereo visual odometry is an evolving field that leverages the advantages of stereo event cameras, which provide asynchronous data from two synchronized event sensors, allowing for 3D depth perception in challenging environments with high-speed motion or varying lighting conditions. The early works on Event-based stereo VO were published in 2018. Henri Rebecq et al. [32] introduced EMVS (Event-based Multi-View Stereo), which enabled 3D reconstruction in real-time using event cameras. This method capitalized on the asynchronous nature of events to compute depth information by fusing event data from multiple viewpoints, enabling accurate and efficient 3D reconstruction. In the same year, a semi-dense 3D reconstruction method using stereo event cameras was proposed [33]. Their approach utilized stereo event cameras to reconstruct semi-dense 3D maps, significantly improving the accuracy and robustness of visual odometry systems, even in dynamic environments. In [34], the authors presented the Multivehicle Stereo Event Camera Dataset, specifically designed for 3D perception. This dataset provided benchmark data for stereo event-based odometry, enabling further research in event-based 3D perception. By providing a large-scale, real-world dataset, this work facilitated the evaluation of stereo event camera systems in various challenging scenarios.
@@
In 2021, Zhou et al. [35] extended their prior work in [33] by introducing a complete event-based stereo visual odometry system. Their method combined stereo event cameras and a robust odometry pipeline to estimate camera motion in real time. By leveraging the high temporal resolution of event cameras, this system was able to track fast-moving objects and scenes with challenging lighting conditions. The feature-based method can also be utilized on event-based stereo visual odometry [36]. Their approach detected and tracked features directly from stereo event data, using them to estimate motion.
@@
Recently, research in stereo event-based visual-inertial odometry (VIO) has gained attention, as it enhances motion estimation by fusing inertial measurements with visual data to improve accuracy and robustness, particularly in conditions where visual information alone may be unreliable. In 2023, Kunfeng Wang, Kaichun Zhao, and Zheng You developed a stereo event-based VIO system, which incorporated both stereo event cameras and inertial measurements to improve the accuracy of pose estimation [37]. Further research into stereo event-based VIO led to the work of Peiyu Chen et al. [38]. In parallel, Liu et al. also presented an event-based stereo visual-inertial odometry system [39]. Their methods demonstrated superior performance in both accuracy and robustness when compared to frame-based VIO systems, particularly in scenarios involving rapid motion.
@@
Most recent works by Junkai Niu et al. have continued to push the boundaries of event-based stereo VO. Their IMU-aided event-based stereo visual odometry system[40] and ESVO2 [41] incorporate direct visual-inertial odometry with stereo event cameras, significantly improving real-time pose estimation accuracy.
@@
III-E Fusion with LiDARs

### v3

@@ Abstract
Event cameras, inspired by biological vision, are asynchronous sensors that detect changes in brightness. They offer notable advantages in environments characterized by high-speed motion, low lighting, or wide dynamic range. These distinctive properties render event cameras particularly effective for sensor fusion in robotics and computer vision, especially in enhancing traditional visual or LiDAR-inertial odometry. Conventional frame-based cameras suffer from limitations such as motion blur and drift, which can be mitigated by the continuous, low-latency data provided by event cameras. Similarly, LiDAR-based odometry encounters challenges related to the loss of geometric information in environments such as corridors. To address these limitations, unlike the existing event camera-related surveys, this survey presents a comprehensive overview of recent advancements in event-based sensor fusion for odometry applications particularly investigating fusion strategies that incorporate frame-based cameras, inertial measurement units, and LiDAR. The survey critically assesses the contributions of these fusion methods to improving odometry performance in complex environments, while highlighting key applications, and discussing the strengths, limitations, and unresolved challenges. Additionally, it offers insights into potential future research directions to advance event-based sensor fusion for next-generation odometry applications.
@@ I Introduction
In the context of robotics and computer vision, sensor fusion involving event cameras has garnered considerable attention, particularly for enhancing visual odometry (VO) or visual-inertial odometry (VIO). Event-based sensor fusion leverages the complementary strengths of event cameras and other sensors (e.g., frame-based cameras, inertial measurement units (IMUs), LiDAR) to improve robustness and accuracy in harsh environments with high-speed needs. Regarding odometry, LiDAR-based systems generally offer greater accuracy and robustness to varying lighting conditions compared to VO or VIO. However, in environments where geometric information is homogenous, such as long corridors, LiDAR odometry is prone to drift [4]. Event cameras, with their ability to capture dynamic scene changes, have the potential to mitigate this issue by providing complementary information that enhances the overall accuracy and robustness of odometry in such challenging scenarios. Fig. 1 illustrates the varying representations of a long corridor as captured by different sensors, including an RGB camera, LiDAR, and an event camera. It is evident from the LiDAR data that the corridor is not clearly detected; however, by integrating the data from the RGB camera and the event camera, this limitation can be mitigated.
@@
The remainder of this paper is structured as follows: Section II provides an overview and fundamentals of event camera technology. Subsequently, Section III discusses of how sensor fusion is used in robotic odometry. Next, Section IV explores the key trends in event camera-based sensor fusion tasks particularly on enhancing odometry by integrating event cameras with other sensors and introducing the most recent datasets relevant to event cameras for odometry. Finally, Section V concludes the paper, offering insights into potential future research directions.
II Event Camera Technology
In robotics and computer vision, event cameras offer several advantages over traditional frame-based cameras, making them particularly suitable for these specific applications [3, 6]. Currently, there are several manufacturers of event cameras, each offering different key features. Table I lists these manufacturers along with their main products and characteristics.
Event cameras capture changes in brightness asynchronously for each pixel. Unlike traditional frame-based cameras that capture intensity images at fixed intervals, event cameras detect changes in the logarithmic brightness of individual pixels independently, producing events whenever a change exceeds a predefined threshold [6]. Specifically, an event at a pixel is triggered if the logarithmic intensity $\Delta L(x,y,t)$ changes since the last event surpasses the contrast threshold:
$\Delta L(x,y,t)=\log I(x,y,t)-\log I(x,y,t_{\text{last}})=p\cdot C$
where $I(x,y,t)$ is the intensity at pixel $(x,y)$ at time $t$; $t_{\text{last}}$ is the time of the last event at pixel $(x,y)$; $p\in\{+1,-1\}$ represents the polarity of the event; $C$ is the contrast threshold determining the camera’s sensitivity to brightness changes.
@@ II-A Comparison with Traditional Cameras
Event cameras capture changes in brightness asynchronously for each pixel. Unlike traditional frame-based cameras that capture intensity images at fixed frame rates, event cameras detect changes in the logarithmic brightness of individual pixels independently, producing events whenever a change exceeds a predefined threshold [6]. Specifically, an event at a pixel is triggered if the logarithmic intensity changes since the last event surpasses the contrast threshold. This difference leads to several key advantages:
@@
III Sensor Fusion for Odometry in Robotics
Traditional VO methods can suffer from challenges such as motion blur, occlusion, and sensitivity to lighting variations. Event cameras, with their high temporal resolution and ability to capture asynchronous changes, offer a promising solution to address these challenges by providing complementary information that enhances the robustness and accuracy of VO. Using sensor fusion, we can integrate the data from these and other sensors, such as visual sensors, IMUs, LiDARs and radars, to achieve more reliable and accurate measurements.
These sensors are used in several studies to enhance robustness of odometry estimation for robotics. Fast-LIO [10][11] combined LiDAR and IMU sensors to achieve odometry, where LiDAR points were extracted as planar and edge features. An iterative Kalman filter was used to fuse these features with IMU measurements, providing accurate state estimation. Building on this, COIN-LIO [12] introduced an improvement by incorporating photometric information from LiDAR intensity data. The system projected LiDAR intensity values into an image format and applied filtering to enhance brightness consistency. The key features in these intensity images were selected based on their complementarity to the geometric features of the LiDAR point cloud, specifically targeting areas with limited geometric information.
@@ III Sensor Fusion for Odometry in Robotics
IV Event-based Sensor Fusion for Odometry
IV-A Event-Camera Only
@@
Since an event camera only detects the change in brightness, it provides less features for localization than frame cameras; therefore, the number of studies using only an event camera for odometry is small. The first event-based pose estimation work was presented by Weikersdorfer and Conradt [16] in 2012. This work used only an event camera and was extended to an event-based visual SLAM system in 2013 [17].
IV-B Fusion with Frame-based Camera
@@
Combining event cameras with frame-based cameras allows for high temporal resolution with more structured scene information. Liu et al. [18] presented an approach to combine both sensor modalities for object detection and tracking. In [19], Pan et al. proposed a method to reconstruct high frame-rate, sharp videos from a single blurry frame and event data using the event-based double integral model. The fusion of asynchronous event data with frame-based data reduces blur and improves visual quality, showcasing the complementary strengths of both modalities. Fusing an event camera with a frame camera can overcome the limitation of each separate sensor and estimate the stereo depth. For example, Wang et al. [20] enabled stereo depth estimation by leveraging both event-based edge information and frame-based intensity data, offering improved disparity estimation compared to state-of-the-art stereo matching algorithms.
@@
The fusion of an event camera and a frame camera can increase the robustness and accuracy of VO. In 2014, Censi and Scaramuzza proposed a low-latency event-based VO system [21], which was a significant advancement in reducing the latency associated with traditional frame-based methods. This system utilized asynchronous event data to compute odometry in real time, paving the way for faster and more efficient pose estimation in dynamic environments. Building upon this, Rebecq et al. [22] introduced low-latency VO using event-based feature tracks. This method tracked feature points based on event data, achieving reliable and low-latency odometry even in conditions where traditional frame-based methods would struggle, such as in fast motion or poor lighting. The same group developed a 6-degree-of-freedom (DoF) event-based tracking and mapping system, which relied on geometric principles to perform both parallel tracking and mapping in real time [23]. This work was instrumental in showcasing the potential of event cameras for performing full 3D tracking and mapping tasks, marking another milestone in event-based VO.
@@
In 2022, Hidalgo-Carrió et al. proposed event-aided direct sparse odometry (DSO), where event data was combined with traditional DSO to improve the robustness of VO in environments with challenging lighting or rapid motion changes [24]. This method demonstrated the fusion of event-based and frame-based approaches in improving VO performance.
IV-C Fusion with IMU
@@
IMUs provide inertial data that can be fused with events to enhance motion estimation, particularly in high-speed scenarios. IMUs can provide estimations of velocity and orientation, aiding visual odometry by reducing drift and improving robustness [25]. In [26], Mohamed et al. explored the use of event data and IMU fusion to achieve enhanced VO performance in dynamic environments. The study proposed methods to leverage the event camera’s high temporal resolution with the IMU’s robust inertial data to accurately estimate motion, even in highly dynamic scenarios. Chamorro et al. proposed new methods for integrating event data and IMU readings to achieve ultra-fast camera pose estimates [27]. They introduced two novel fusion schemes that combine constant velocity and constant acceleration prediction models with 10 kHz ultra-fast event-based updates and slower 1 kHz IMU updates. The findings demonstrate superior performance at a throughput that is 100 times faster than state-of-the-art.
@@
Integrating an IMU into VO systems significantly enhances their performance by providing additional, reliable motion information that compensates for the limitations of visual data alone. In 2017, Zhu et al. extended the capabilities of event-based odometry by integrating inertial measurements in their work on event-based visual-inertial odometry [28]. The combination of event cameras and inertial sensors allowed for robust pose estimation even in challenging scenarios, such as high-speed motion or complex dynamic scenes. In 2018, Mueggler and his team presented a continuous-time visual-inertial odometry approach for event cameras, further enhancing the accuracy and robustness of event-based odometry systems [29]. Their work highlighted the importance of fusing temporal information from both event data and inertial sensors in a continuous-time framework.
@@
Mahlknecht et al. [30] explored event camera-based odometry for planetary robots in 2022, showcasing the potential of event cameras in extraterrestrial environments where traditional cameras may struggle due to harsh conditions. In 2023, Lee et al. [31] introduced a hybrid approach combining event and frame-based visual-inertial odometry with adaptive filtering techniques based on 8-DoF warping uncertainty. This method enhanced the odometry system’s resilience to noise and dynamic changes, further improving the robustness and adaptability of event-based systems in various application settings.
IV-D Event-based Stereo Visual Odometry
@@
Event-based stereo visual odometry is an evolving field that leverages the advantages of stereo event cameras, which provide asynchronous data from two synchronized event sensors, allowing for 3D depth perception in challenging environments with high-speed motion or varying lighting conditions. The early works on Event-based stereo VO were published in 2018. Rebecq et al. [32] introduced Event-based Multi-View Stereo VO, which enabled 3D reconstruction in real-time using event cameras. This method capitalized on the asynchronous nature of events to compute depth information by fusing event data from multiple viewpoints, enabling accurate and efficient 3D reconstruction. In the same year, a semi-dense 3D reconstruction method using stereo event cameras was proposed [33]. Their approach utilized stereo event cameras to reconstruct semi-dense 3D maps, significantly improving the accuracy and robustness of visual odometry systems, even in dynamic environments.
@@
In 2021, Zhou et al. [34] extended their prior work in [33] by introducing a complete event-based stereo visual odometry system. Their method combined stereo event cameras and a robust odometry pipeline to estimate camera motion in real time. By leveraging the high temporal resolution of event cameras, this system was able to track fast-moving objects and scenes with challenging lighting conditions. The feature-based method can also be utilized on event-based stereo visual odometry [35, 36]. Their approach detected and tracked features directly from stereo event data, using them to estimate motion.
@@
Recently, research in stereo event-based VIO has gained attention, as it enhances motion estimation by fusing inertial measurements with visual data to improve accuracy and robustness, particularly in conditions where visual information alone may be unreliable. In 2023, Wang et al. developed a stereo event-based VIO system, which incorporated both stereo event cameras and inertial measurements to improve the accuracy of pose estimation [37]. Further research into stereo event-based VIO led to the work of Chen et al. [38]. In parallel, Liu et al. also presented an event-based stereo visual-inertial odometry system [39]. Their methods demonstrated superior performance in both accuracy and robustness when compared to frame-based VIO systems, particularly in scenarios involving rapid motion.
@@
Most recent works by Niu et al. have continued to push the boundaries of event-based stereo VO. Their IMU-aided event-based stereo visual odometry system[40] and ESVO2 [41] incorporate direct visual-inertial odometry with stereo event cameras, significantly improving real-time pose estimation accuracy.
@@
2023 UNIZG-FER LAMOR [36] Indoor handheld & Outdoor Driving Stereo Event Camera, LiDAR, IMU, GNSS
@@
2018 MVSEC [46] Handheld, UAV & Driving Stereo Event Camera, LiDAR, VI-Sensor, IMU, GNSS
@@
IV-E Fusion with LiDARs